# Optical pumping of alkali-metal vapor with hyperfine-resolved buffer gas pressure

### 06 Mar 2026 v1

@@ III.3.2 Steady state population
with the kernel $\mathcal{G}_{\mathrm{pop}}$
$\mathcal{G}_{\mathrm{pop}}=$ $\Gamma_{\mathrm{SD}}\left[A_{\mathrm{SD}}\right]_{0,0}$
$+R_{\mathrm{op}}\left(\left[A_{\mathrm{SD}}\right]_{0,0}-\frac{s}{2}\left[A^{(z)}_{\mathrm{SE}}\right]_{0,0}\right)\mathcal{Q}^{\{0\}}_{\mathcal{L}}$
$+\Gamma_{\mathrm{ex}}\left(\left[A_{\mathrm{SD}}\right]_{0,0}-(S_{z}|\rho_{0})\left[A^{(z)}_{\mathrm{SE}}\right]_{0,0}\right),$ (45)
where the average electron spin $(S_{z}|\rho_{0})=\mathrm{Tr}\left[S_{z}\rho_{0}\right]$. We explore this equation as a Markov chain [16] of classical probability, and particularly emphasis on its unique steady state $\left|\rho_{\mathrm{st}}\right)=\left|\rho_{0}(t\to+\infty)\right)$, which is most relevant in practice. Here, $\left|\rho_{\mathrm{st}}\right)$ is given by the null space of $\mathcal{G}_{\mathrm{pop}}$. We found that the unstretched state with $m\in[-b,\,b]$ generally satisfies
$\frac{(am;am\left|\rho_{\mathrm{st}}\right)}{(bm;bm\left|\rho_{\mathrm{st}}\right)}$ $=\frac{\Gamma_{\mathrm{ex}}+\Gamma_{\mathrm{SD}}+R_{\mathrm{op}}{Q}_{b}}{\Gamma_{\mathrm{ex}}+\Gamma_{\mathrm{SD}}+R_{\mathrm{op}}{Q}_{a}}.$ (46)
This directly illustrates the effective pumping effect between the $a$ and $b$ multiplets, which vanishes in the HP limit. We next turn to specific cases corresponding to different mean photon spin $s$.
For circularly polarized photons such as $s=1$, the steady state no longer satisfies the detailed balance, namely spin temperature distribution (STD) [11, 1], which can be obtained only with ${Q}_{{F}}=1$. The numerical solution in Fig. 4(a) describes its deviation from STD. Also in contrast to the HP limit, we further found that the electron spin polarization $P$, defined as
$P\equiv 2\langle S_{z}\rangle=2(S_{z}|\rho_{\mathrm{st}}),$ (47)
becomes dependent on the spin exchange rate $\Gamma_{\mathrm{ex}}$, as shown in Fig. 4(b), This is because when the rapidly oscillating hyperfine coherence decays, spin exchange collision no longer conserves spin for states beyond STD.
[…]
$\left(Fm_{F};Fm_{F}\left|\rho^{(\mathrm{STD})}_{\mathrm{st}}\right)\right.=\frac{P(1+P)^{\frac{[I]}{2+m_{F}}}(1-P)^{\frac{[I]}{2-m_{F}}}}{(1+P)^{[I]}-(1-P)^{[I]}}.$ (48)
In Sec. IV, we will give $p_{F}$, $\langle S_{z}\rangle_{F}$ and $C_{F}$ in Eqs. (41), (42) and (43) in the STD limit, which enables renewed absorption cross section $\sigma$ and polarization $P$.
III.3.3 Weak driving magnetic resonance
A practical application of longitudinal pumping is atomic magnetometers based on magnetic resonance, which lies in the population imbalance among $m_{F}$ sublevels created by circularly polarized photons, with $s=1$ for simplicity. This essentially enables the transverse RF field to drive Zeeman coherences [1]. In our previous work for unresolved Zeeman resonance [22], we considered an radio-frequency (RF) driving field along the $\hat{x}$ axis
$H_{\mathrm{drv}}=\Omega_{\mathrm{R}}S_{x}\cos(\omega_{\mathrm{rf}}t)$ (49)
whose oscillation frequency $\omega_{\mathrm{rf}}$ far exceeds the Larmor frequency $\Omega_{\mathrm{L}}$, i.e., the linear Zeeman splitting. We assumed that the Rabi frequency $\Omega_{\mathrm{R}}$ is weak enough to keep $\left|{\rho}_{\mathrm{st}}\right)$ unchanged and only 1st-order Zeeman coherence $\left|\tilde{\rho}_{1}\right)$ generated for steady state. Therefore, a framework in the rotating RF frame is established for $\left|\tilde{\rho}_{1}\right)$
$\left|\tilde{\rho}_{1}\right)=$ $-\frac{i\Omega_{\mathrm{R}}}{2}\mathcal{G}^{-1}_{\mathrm{\mathrm{WDA}}}\big[S^{\copyright}_{x}\big]_{1,0}\left|{\rho}_{\mathrm{st}}\right),$ (50)
where $\big[S^{\copyright}_{x}\big]_{1,0}$ generates $\left|\tilde{\rho}_{1}\right)$ from polarized atomic population $\left|{\rho}_{0}\right)$, and $\mathcal{G}_{\mathrm{\mathrm{WDA}}}$ is a total relaxation superoperator of $\left|\tilde{\rho}_{1}\right)$. Following this framework, the optical pumping contribution should be replaced by $\left[A_{\mathrm{op}}\right]_{1,1}$, whose real part related to decoherence is shown in Fig. 3, and the imaginary part leading to frequency shifts is
$\mathrm{Im}\left[\mathcal{A}_{\mathrm{op}}\right]_{1,\,1}$ $=\frac{\Delta_{a}Q_{a}}{[I]\Gamma_{\mathrm{brd}}}\mathcal{P}^{\{1\}}_{a}+\frac{\Delta_{a}Q_{b}}{[I]\Gamma_{\mathrm{brd}}}\mathcal{P}^{\{1\}}_{b}$. (51)
Meanwhile, all other relaxation terms in $\mathcal{G}_{\mathrm{\mathrm{WDA}}}$ remain formally unchanged.
In typical magnetometers, transverse spin component is measured via the Faraday rotation effect [10, 4]. Accordingly, the unresolved Zeeman resonance can be well described by a complex response amplitude [22]
@@ III.3.3 Weak driving magnetic resonance
$\mathrm{S}=(S_{x}\left|\tilde{\rho}_{1}\right)\approx$ $\frac{\Omega_{\mathrm{R}}}{2}\frac{W_{\mathrm{}}}{\Delta_{\mathrm{rf}}+i\Gamma_{2}}$ (52)
where $\Delta_{\mathrm{rf}}=\omega_{\mathrm{rf}}-\Omega_{\mathrm{L}}-\delta\Omega_{\mathrm{L}}$ is the effective RF detuning, with $\delta\Omega_{\mathrm{L}}$ being the optical shift, i.e. light-induced AC-Stark shift to $\Omega_{\mathrm{L}}$. The linewidth $\Gamma_{2}$ is the lowest eigenvalue of $\mathcal{G}_{\mathrm{\mathrm{WDA}}}$ with its weight factor $W$, which strongly dominates the exact $\mathrm{S}$. This eigenvalue exists in the a-subspace, which renders $\delta\Omega_{\mathrm{L}}$ determined by the corresponding frequency shift in Eq. (51)
$\delta\Omega_{\mathrm{L}}$ $\approx\frac{1}{[I]}\frac{\Phi A_{0}\Delta_{a}}{\Delta^{2}_{a}+\Gamma^{2}_{\mathrm{brd}}}.$ (53)
As observable signals, the in-phase and out-of-phase quadratures are
$X=\mathrm{Re}\left[\mathrm{S}\right]=\frac{\Omega_{\mathrm{R}}}{2}\frac{W_{\mathrm{}}\Delta_{\mathrm{rf}}}{\Delta^{2}_{\mathrm{rf}}+\Gamma^{2}_{2}},$ (54)
$Y=-\mathrm{Im}\left[\mathrm{S}\right]=\frac{\Omega_{\mathrm{R}}}{2}\frac{W_{\mathrm{}}\Gamma_{2}}{\Delta^{2}_{\mathrm{rf}}+\Gamma^{2}_{2}}.$ (55)
The amplitude $Y_{\mathrm{max}}=\Omega_{\mathrm{R}}W/\Gamma_{2}$ with $\Delta_{\mathrm{rf}}=0$ describes the maximum response of atomic magnetometers.
In Sec. IV.3, observables associated with atomic magnetometer performance, $\Gamma_{2}$ and $Y_{\mathrm{max}}$, will be investigated in detail to demonstrate how the magnetic resonance is enhanced at specific laser detuning. The framework in the QHP regime can be straightforwardly extended to the study of resolved Zeeman resonance [1, 2, 14], with only modifications accounting for the contribution of optical pumping with the F=a,bF=a,\,b multiplets resolved.
@@ IV.1 Absorption cross section
For optically pumped atoms in the QHP regime, we utilize the description of light absorption in Sec. III.3.1 and calculate $\sigma_{\mathrm{lin}}$ and $\sigma_{\mathrm{cir}}$.
For linearly polarized pumping with $s=0$, the atomic spin cannot be polarized and Eq. (44) exactly result in detailed balance [16]. Thus, we can analytically get $\left|\rho_{\mathrm{st}}\right)$ by Eq. (46).
$\left|\rho^{(s=0)}_{\mathrm{st}}\right)$ $=\dfrac{1}{N_{0}}\sum_{F}\sum_{m_{F}}{\left(1+\dfrac{R_{\mathrm{op}}Q_{F^{\prime}}}{\Gamma_{\mathrm{SD}}+\Gamma_{\mathrm{ex}}}\right)\Big|Fm_{F};Fm_{F}\Big)},$ (56)
with $F^{\prime}\neq F$ denoting the other hyperfine multiplet distinct from $F$, and $N_{0}=2[I]+\big([a]Q_{b}+[b]Q_{a}\big)R_{\mathrm{op}}$ being the normalization factor. Due to unresolved $m_{F}$ sublevels within the $F$ multiplet, an effective three-level model featured in the QHP regime naturally emerges—comprising $F=a$ and $F=b$ together with the ES—capturing the optical excitation and relaxation pathways. For the steady state given by Eq. (56), $R_{\mathrm{op}}Q_{F}$ plays the role of an effective transfer between the GS hyperfine multiplets, while $\Gamma_{\mathrm{ex}}+\Gamma_{\mathrm{SD}}$ characterizes the relaxation-induced population exchange between the two hyperfine multiplets, leading to a distribution governed by their degeneracies. Populations in the hyperfine multiplets $F=a$ and $F=b$ contribute to the light absorption differently, and $\sigma_{\mathrm{lin}}$ can be expressed as
$\sigma_{\mathrm{lin}}$ $=C^{(\mathrm{lin})}_{a}\,\frac{A_{0}\Gamma_{\mathrm{brd}}}{\Delta_{a}^{2}+\Gamma_{\mathrm{brd}}^{2}}+C^{(\mathrm{lin})}_{b}\,\frac{A_{0}\Gamma_{\mathrm{brd}}}{\Delta_{b}^{2}+\Gamma_{\mathrm{brd}}^{2}},$ (57)
where the coefficients $C^{(\mathrm{lin})}_{a}$ and $C^{(\mathrm{lin})}_{b}$ take the same form $F^{\prime}\neq F$ same as in Eq.(56)]
$C^{(\mathrm{lin})}_{F}$ $=\dfrac{\Gamma_{\mathrm{SD}}+\Gamma_{\mathrm{ex}}+R_{\mathrm{op}}Q_{F^{\prime}}}{2[I](\Gamma_{\mathrm{SD}}+\Gamma_{\mathrm{ex}})+\big([a]Q_{b}+[b]Q_{a}\big)R_{\mathrm{op}}},$ (58)

### 14 Apr 2026 v2

@@ III.3.2 Population and polarization
with the superoperator $\mathcal{G}_{\mathrm{pop}}$
$\mathcal{G}_{\mathrm{pop}}=$ $\Gamma_{\mathrm{SD}}\left[\mathcal{A}_{\mathrm{SD}}\right]_{0,0}+R_{\mathrm{op}}\left[\mathcal{A}_{\mathrm{op}}\right]_{0,0}$
$+\Gamma_{\mathrm{ex}}\left(\left[\mathcal{A}_{\mathrm{SD}}\right]_{0,0}-\langle S_{z}\rangle\left[\mathcal{A}^{(z)}_{\mathrm{SE}}\right]_{0,0}\right),$ (38)
where $\langle S_{z}\rangle=(S_{z}|\rho_{0})$ is the expectation value of the electron spin along the $z$ direction.
The steady-state solution, $\left|\rho_{\mathrm{ss}}\right)=\left|\rho_{0}(t\to+\infty)\right)$, is obtained by evaluating the null space of $\mathcal{G}_{\mathrm{pop}}$. Within the QHP regime, the population ratio, $\eta$, between the $|F=a,m_{F}\rangle$ and $|F=b,m_{F}\rangle$ states for a given $m_{F}$ ($|m_{F}|\neq a$) is given by
$\eta=\frac{(a,m_{F};a,m_{F}\left|\rho_{\mathrm{ss}}\right)}{(b,m_{F};b,m_{F}\left|\rho_{\mathrm{ss}}\right)}=\frac{\Gamma_{\mathrm{ex}}+\Gamma_{\mathrm{SD}}+R_{\mathrm{op}}{Q}_{b}}{\Gamma_{\mathrm{ex}}+\Gamma_{\mathrm{SD}}+R_{\mathrm{op}}{Q}_{a}}.$ (39)
This ratio is independent of $m_{F}$. In the conventional HP limit, the absorption efficiencies become equivalent ($Q_{a}=Q_{b}$), driving the ratio to $\eta\rightarrow 1$. This limit seamlessly recovers the widely recognized spin-temperature distribution (STD) [11, 1]. However, in the scenario where $Q_{a}\neq Q_{b}$, the system diverges from this standard behavior. Under these conditions, the asymmetric absorption efficiencies parameterized by $Q_{F}$ induce an effective population transfer between the $F=a$ and $F=b$ multiplets, a phenomenon that dictates the core dynamics explored in the subsequent analysis.
For linearly polarized pumping with effective photon spin $s=0$, the population distribution is exactly solvable. Since no angular momentum is transferred from the incident field to the atomic system, the $m_{F}$-pumping process is precluding. Consequently, these magnetic sublevels remain uniformly populated, and the generation of spin polarization is strictly forbidden (i.e., $\langle S_{z}\rangle_{F}=\langle S_{z}\rangle\equiv 0$). The system dynamics therefore reduce to those of an effective two-level system comprising the hyperfine multiplets $F=a$ and $F=b$. Under these conditions, the analytical steady-state solution, $\left|\rho_{\mathrm{lin}}\right)$, dictated by the principle of detailed balance [16], is given by
$\left|\rho_{\mathrm{lin}}\right)$ $=\sum_{F,m_{F}}\frac{p_{F}}{2F+1}\left|F,m_{F};F,m_{F}\right),$ (40)
$p_{F}$ $=\frac{x_{F}\left(\Gamma_{\mathrm{SD}}+\Gamma_{\mathrm{ex}}+R_{\mathrm{op}}Q_{F^{\prime}\neq F}\right)}{\Gamma_{\mathrm{SD}}+\Gamma_{\mathrm{ex}}+R_{\mathrm{op}}(x_{a}Q_{b}+x_{b}Q_{a})},$ (41)
where $\Gamma_{\mathrm{ex}}+\Gamma_{\mathrm{SD}}$ denotes the relaxation-induced population exchange rate between the multiplets, acting as a restorative mechanism that drives the system toward the equilibrium. Conversely, the term $R_{\mathrm{op}}Q_{F^{\prime}\neq F}$ characterizes the effective population transfer from the $F^{\prime}$ to the $F$ multiplet, creating the unbalanced population between the two manifolds under the linearly polarized pumping.
Figure 4: (a) Representative steady-state populations under slow spin-exchange conditions.. (b) Total population of the $F=a$ multiplet as a function of the normalized pumping rate $R_{\mathrm{op}}/\Gamma_{\mathrm{SD}}$, with curves denoting different $Q_{F}$ configurations. (c) The 1-norm distance $d$ [see Eq. (42)] between the steady-state population and an STD profile of the equivalent polarization, fixing $R_{\mathrm{op}}/\Gamma_{\mathrm{SD}}=1$. (d) Spin polarization $P$ versus $\Gamma_{\mathrm{ex}}/\Gamma_{\mathrm{SD}}$. Color coding for (c) and (d) is consistent with (b). Results are obtained for $I=3/2$ and $K=s=1$.
Under circularly polarized pumping ($s=\pm 1$), the system dynamics become increasingly complex due to the concurrent action of $F$-pumping and $m_{F}$-pumping. Because these coupled processes are governed by the non-linear Eq. (37), which lacks an analytical solution, the steady-state distributions are solved via numerical simulation.
Specifically, in the standard HP limit, solely $m_{F}$-pumping dictates the dynamics; this mechanism drives the population toward states with higher $m_{F}$ values while maintaining the population balance between the $F=a$ and $F=b$ manifolds for a given $m_{F}$. Conversely, within the QHP regime, the emergence of $F$-pumping breaks this multiplet balance. As illustrated in Fig. 4(a), when $Q_{a}>Q_{b}$, the disproportionate absorption within the $F=a$ manifold pumps the population into the $F=b$ levels (red bars). An analogous population redistribution occurs under the $Q_{b}>Q_{a}$ condition (blue bars). Figure 4(b) elucidates the dependence of this impact on the normalized pumping intensity, parameterized as $R_{\mathrm{op}}/\Gamma_{\mathrm{SD}}$. This redistribution mechanism exerts a negligible influence when approaching either the weak-pumping limit ($R_{\mathrm{op}}\ll\Gamma_{\mathrm{SD}}$) of a nearly uniform distribution or the strong-pumping limit ($R_{\mathrm{op}}\gg\Gamma_{\mathrm{SD}}$) of stretched-state saturation. In contrast, the intermediate domain obviously manifests the $F$-pumping dynamics driven by $Q_{a}\neq Q_{b}$, where, for instance, the notably suppressed red curve (compared with the dashed line) illustrates the net transfer from the $F=a$ to the $F=b$ levels.
The $F$-pumping-induced imbalance in the QHP regime drives the system away from the conventional STD characteristic of the HP limit. Specifically, the steady-state density matrix can no longer be parameterized by a single temperature factor $\beta(P)$ in the form $|\rho_{\mathrm{STD}})\propto\exp[\beta(P)F_{z}]$ [11, 1], with $F_{z}=S_{z}+I_{z}$ being the $z$-component of the total angular momentum operator ($I_{z}$ the nuclear spin counterpart), and $P=2(S_{z}|\rho_{\mathrm{ss}})$ being the steady-state spin polarization . To quantify this deviation, we define a 1-norm distance metric
$d=\|\rho_{\mathrm{ss}}-\rho_{\mathrm{STD}}\|_{1}$ (42)
which compares the QHP steady-state distribution to an ideal STD possessing identical spin polarization $P$. Figure. 4(c) plots this distance as a function of $\Gamma_{\mathrm{ex}}/\Gamma_{\mathrm{SD}}$—the ratio of the spin-exchange rate to the spin-destruction rate—at a fixed optical pumping rate $R_{\mathrm{op}}$. The distance $d$ decreases monotonically as the ratio increases. This demonstrates that rapid spin-exchange collisions provide a dominant restorative mechanism, effectively thermalizing the system to counteract the $F$-pumping induced by asymmetric absorption. The high validity of the spin-temperature approximation under the strong spin-exchange condition ($\Gamma_{\mathrm{ex}}/\Gamma_{\mathrm{SD}}\gg 1$) establishes a critical theoretical foundation for utilizing the STD to simplify the physical models of various practical applications, such as SERF magnetometers, even when the HP condition is not necessarily fully satisfied.
Although the STD serves as a valid approximation at high spin-exchange rate, the joint contribution of $F$- and $m_{F}$-pumping under QHP condition gives rise to a spin polarization different to that obtain in the HP condition. In the HP limit with $Q_{a}=Q_{b}$ [see the black dashed line in Fig. 4(c)], the spin polarization is
$P_{\mathrm{HP}}=\frac{KsR_{\mathrm{op}}}{R_{\mathrm{op}}+\Gamma_{\mathrm{SD}}}$ (43)
In contrast, when $Q_{a}\neq Q_{b}$, the polarization $P_{\mathrm{QHP}}$ converges to a stable value at high spin-exchange rate. The asymptotic polarization under QHP conditions (blue and red dashed lines) is sensitive to the asymmetric absorption strengths ($Q_{a}$ and $Q_{b}$). Consequently, within the QHP framework, the spin polarization $P$ under the STD approximation cannot be obtained trivially; and must be reshaped by the imbalance $Q_{a}\neq Q_{b}$. Detailed results and their corresponding observable effects will be presented in Section IV.
III.3.3 Intrinsic magnetic resonance
The population imbalance across the $m_{F}$ sublevels enables Zeeman magnetic resonance under a radio-frequency (RF) driving field oscillating at $\omega_{\mathrm{rf}}$. To streamline the theoretical analysis, we conceptualize the framework established in our previous work [22] under the notion of intrinsic magnetic resonance. This intrinsic regime describes the 1st-order Zeeman coherence generated by an infinitesimally weak driving field, operating under a weak static magnetic field where the non-linear Zeeman effect is negligible, so that the spin precession frequencies degenerate into a singular Larmor frequency $\omega_{\mathrm{L}}$. In this case, the resonance dynamics are captured by the generated 1st-order Zeeman coherence $\left|\tilde{\rho}_{1}\right)$. In the RF rotating frame, the coherencce is governed by $\left|\tilde{\rho}_{1}\right)\sim\Omega_{\mathrm{R}}\mathcal{G}^{-1}_{\mathrm{\mathrm{WDA}}}\big[S^{\copyright}_{x}\big]_{1,0}\left|{\rho}_{\mathrm{st}}\right),$ where $\Omega_{\mathrm{R}}$ is the weak Rabi frequency, $\big[S^{\copyright}_{x}\big]_{1,0}$ dictates the coherence generation from polarized atoms, and $\mathcal{G}_{\mathrm{WDA}}$ is the total relaxation superoperator under the weak driving approximation. When the resonance is measured by observing the transverse spin component via Faraday rotation [10, 4], and the signal is characterized by the complex amplitude
@@ III.3.3 Intrinsic magnetic resonance
$R=(S_{x}\left|\tilde{\rho}_{1}\right)\approx$ $\frac{\Omega_{\mathrm{R}}}{2}\frac{W_{\mathrm{}}}{\Delta_{\mathrm{rf}}+i\Gamma_{2}}$ (44)
Here, $W$ represents the dominant Zeeman coherence weight, and $\Delta_{\mathrm{rf}}=\omega_{\mathrm{rf}}-\omega_{\mathrm{L}}-\delta\omega_{\mathrm{L}}$ defines the effective RF detuning, taking into account the optical shift $\delta\omega_{\mathrm{L}}$. The linewidth $\Gamma_{2}$ is obtained by the minimal eigenvalue of $\mathcal{G}_{\mathrm{WDA}}$, and the optical shift originates from the value of $\mathrm{Im}[\mathcal{A}_{\mathrm{op}}]_{1,1}$ [22]. Experimentally, properties of the resonance directly manifest as the in-phase and out-of-phase quadratures, i.e., $X=\mathrm{Re}[R]$ and $Y=-\mathrm{Im}[R]$, respectively.
Transitioning this framework into the QHP regime necessitates a $Q_{F}$-induced modification. While all other relaxation mechanisms in $\mathcal{G}_{\mathrm{WDA}}$ remain formally unchanged, the optical pumping contribution to $\left|\tilde{\rho}_{1}\right)$ relaxation is replaced by the hyperfine-resolved superoperator $\left[\mathcal{A}_{\mathrm{op}}\right]_{1,1}$ [see Eq. (31) & Fig. 3] to incorporate the $Q_{a}\neq Q_{b}$ absorption asymmetry. Detailed results about the linewidth calculations and the magnetic resonance enhancement under QHP conditions are discussed in Sec. IV.3.
@@ IV.1 Absorption cross section
For optically pumped atoms in the QHP regime, we utilize the description of light absorption in Sec. III.3.1 to calculate the absorption cross-section $\sigma_{\mathrm{lin}}$ and $\sigma_{\mathrm{cir}}$ for linearly- and circularly-polarized laser beams.
For linearly polarized pumping ($s=0$), populations in the hyperfine multiplets $F=a$ and $F=b$ contribute to the light absorption differently, and $\sigma_{\mathrm{lin}}$ is
$\sigma_{\mathrm{lin}}$ $=A\left[p_{a}\mathcal{L}(\Delta_{a})+p_{b}\mathcal{L}(\Delta_{b})\right],$ (45)